OpenStudy (anonymous):

if g(x) = x2 + 3, find g(4).

jimthompson5910 (jim_thompson5910):

g(x) = x^2 + 3 g(4) = 4^2 + 3 g(4) = ???

OpenStudy (anonymous):

so it would be 11 or 8 because the first time i did it i got 19 .

jimthompson5910 (jim_thompson5910):

g(x) = x^2 + 3 g(4) = 4^2 + 3 g(4) = 16 + 3 g(4) = 19

jimthompson5910 (jim_thompson5910):

so you were correct the first time

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np